Consider the reaction: Fe2O3(s) + 2Al(s)Al2O3(s) + 2Fe(s) Using standard absolute entropies at 298K, calculate the entropy change for the system when 2.37 moles of Fe2O3(s) react at standard conditions. S°system = ____J/K
Answer #1

Fez 03.c. +24 609 4103 603 + 2€ (3) As = sp - Sk = { $440, +2(3e)] - [Specs + 2CM)] = [50-986 +2 (27-2)] - [ 90+2C261 327) =